Week of 160905

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Maria Alandes (chair and minutes), Thomas Javurek (ATLAS), Maarten Litmaath (ALICE), Ignacio Coterillo (DB), Giuseppe Lo Presti and Alessandro Fiorot (Storage), Marian Babik (Networks), Vincent Brillault (Security), Julia Andreeva (WLCG)
  • remote: Andrei Tsaregorodtsev (LHCb), Christoph Wissing (CMS), Elena Corni (CNAF), Vincenzo Spinoso (EGI), Dave Mason (FNAL), Rolf Rumler (IN2P3), Victor Zhiltsov (JINR), Sang-Un Ahn (KISTI), Dmytro Karpenko (NDGF), Onno Zweers (NL-T1), Eygene Ryabinkin (RRC-KI), Pepe Flix (pic), Tiju Idiculla (RAL), Alberto Aimar (Monitoring)

Experiments round table:

  • ATLAS reports (raw view) -
    • Activities:
      • Spillover tasks: testing, new feature of Athena - event forking after first event, some issues with comparison of DQ histos
      • Reprocessing of physics_main date is planed, RAW2ALL test task ongoing http://bigpanda.cern.ch/task/9348904/
      • Protocols;
        • intensive effort to set gsiftp sites in US cloud
        • https/webdav used for deletion at many sites
    • Problems:
      • Winger to EOS stream stalled on Sunday for 30 minutes, no related issues
      • Kibana monitoring issues.
      • AMI service appeared red for two days in our monitoring, it is fixed now.
      • FZK-LCG2 space issues, after deletion it took three days than the space decrease was reported

Maria suggests to open GGUS tickets for the kibana instabilities so that it is possible to follow up the issues.

  • CMS reports (raw view) -
    • CERN EOS instance lost subscription to the xrootd redirector infrastructure: GGUS:123719
    • Some instabilities of the monitoring infrastructure observed during the weekend
      • Occasionally some Kibana pages suffered from missing input data
      • Not really easy to understand
    • Saw some limitations in P5 to Tier-0 data transfers during the weekend
      • Logging rate was rather high
      • Correlation with outage of one Wigner network link not 100% clear
    • Some (permission?) issues with writing of Log files to Castor: GGUS:123701
      • Still under investigation
    • Found some IPv6 related problems in the AAA xrootd federation
      • Experts could mitigate issues by a workaround
    • Tape deletion campaign
      • Deletion requests sent to sites end of last week
      • Requests for CCIN2P3 and PIC were auto-approved (which was not planned)
      • Emergency stop of deletions today
      • Even after weeks of circulating the deletion lists some new exceptions were requested

As ATLAS, Maria suggests to open GGUS tickets for the kibana instabilities so that it is possible to follow up the issues.

Giuseppe reports that he updated the ticket on the log files issues in CASTOR and that the problem was related to wrong permissions on a directory automatically created by CMS.

  • ALICE -
    • NTR

  • LHCb reports (raw view) -
    • Activity
      • Monte Carlo simulation, data reconstruction/stripping and user jobs on the Grid
    • Site Issues
      • T0:
        • T-Systems cloud extension is in draining for the LHCb jobs
      • T1:
        • PIC new HTCondorCE ce13.pic.es is deployed and added to BDII and being added to the LHCb configuration
        • Disk server failure at RAL, 16 files lost

Maria suggests LHCb to inform the right people within CERN IT coordinating the T-systems cloud extension about the fact that LHCb resources could be used by other VOs.

Sites / Services round table:

  • ASGC: NA
  • BNL: NA
  • CNAF: NTR
  • EGI: NTR
  • FNAL: 24h downtime for the whole site on 17th September due to transformer works.
  • GridPP: NA
  • IN2P3: A very minor issue: we are waiting for a reply for the CMS user ticket GGUS:123576 since one week.
  • JINR: Connection to T2_BE_IIHE via LHCONE failed. Checking it.
  • KISTI: NTR
  • KIT: NA
  • NDGF:
    • UmeŚ computing site in downtime 06.09 for software upgrade.
    • Storage in downtime on 08.09 for software upgrade. Data completely unavailable from 08:00 to 09:00 UTC (hopefully much shorter).
  • NL-T1: SARA will be in down time during two weeks in October to migrate to the new data center.
  • RRC-KI-T1: NTR
  • OSG: NA
  • PIC: Downtime on 14th September to upgrade dCache and Network to 20gbps
  • RAL: NTR
  • TRIUMF: NA

  • CERN computing services: NA
  • CERN storage services:
    • EOS updates for CMS and LHCb next week during the LHC Technical Stop: OTG:0032760 and OTG:0032761.
    • Investigating a data loss case due to a client-side race condition (GGUS:123566). Need to contact CMS Phedex developers for following this up.
  • CERN databases: NTR
  • GGUS: NA
  • Monitoring:
  • MW Officer: NTR
  • Networks:
    • GGUS:121687 RAL consistent loss - waiting for router upgrade - plan is to upgrade by end of Sept.
    • MIT inbound throughput - started July 7, was narrowed down to Internet2 to MIT segment, ticket submitted to Internet2.
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r15 < r14 < r13 < r12 < r11 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r15 - 2016-09-05 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback